### v2

# Testing Low-Resource Language Support in LLMs Using Language Proficiency Exams: the Case of Luxembourgish

Abstract
Large Language Models (LLMs) have become an increasingly important tool in research and society at large. While LLMs are regularly used all over the world by experts and lay-people alike, they are predominantly developed with English-speaking users in mind, performing well in English and other wide-spread languages while less-resourced languages such as Luxembourgish are seen as a lower priority. This lack of attention is also reflected in the sparsity of available evaluation tools and datasets. In this study, we investigate the viability of language proficiency exams as such evaluation tools for the Luxembourgish language. We find that large models such as ChatGPT, Claude and DeepSeek-R1 typically achieve high scores, while smaller models show weak performances. We also find that the performances in such language exams can be used to predict performances in other NLP tasks.
1 Introduction
Large Language Models (LLMs) have become increasingly ubiquitous in recent research, with new models being frequently released and updated. They have become similarly omnipresent in the everyday lives of private and professional users, with models being deployed in wide-ranging domains such as finance, medicine, and coding.
While LLMs are primarily catered towards speakers of widespread languages such as English, Spanish, and Chinese, low-resource languages are generally regarded as a lower priority. Indeed, many major LLMs such as the Llama family (Touvron et al., , Touvron et al., 2023, Grattafiori et al., 2024) are trained nearly exclusively on English data, and even LLMs with a focus on non-English languages such as Bloom Le Scao et al. (2023) barely include low-resource languages. As a consequence, numerous mistakes can occur due to cross-lingual transfer between related languages. These mistakes include mistranslated words and partial code-switching, incorrect word order, misgendered nouns, and violation of language-exclusive rules such as the n-rule (Eifeler Regel) in the Luxembourgish language Gilles (2006).
As such, it is important to determine the capabilities of models to understand and generate texts in such languages, and to recognise when they can fail. The benefits of such research are two-fold as it allows (1) to find general weaknesses and pitfalls that can be addressed and mitigated in future LLMs; (2) to make recommendations on which LLM to use for a given scenario in a given low-resource language. The latter is especially important in use cases where an LLM would have to handle sensitive data such as financial or medical information that is not allowed to leave a financial or medical institution due to internal policies and data privacy laws such as the Health Insurance Portability and Accountability Act (HIPAA) Act (1996) or the the General Data Protection Regulation (GDPR) Regulation (2016). In such cases, data must be processed using locally deployed models that do not access the internet.
In this study, we aim to investigate the capabilities of LLMs for Luxembourgish, a low-resource language spoken by nearly 600 000 people world-wide111according to https://cursus.edu/en/23040/luxembourgish-at-its-best, but predominantly in Luxembourg. We want to determine those capabilities from two angles, (a) how well LLMs perform in linguistic proficiency exams used in language institutions; (b) whether their performance in such exams can predict their performance in text generation tasks.
Specifically, we investigate 45 LLMs from a wide array of LLM families that we compare systematically. In addition, we ascertain in which cases LLMs fail and discover trends and common weaknesses using language exams. Finally, we aim to determine whether or not such exams are a useful tool to predict how well an LLM performs in two tasks related to text summarisation. While this study focuses on the Luxembourgish language, we believe that our methodology can be generalised to other low-resource languages as well.
In summary, our contributions are as follows:
  1. (a)
    A systematic comparison of 45 popular LLMs with regard to their linguistic proficiency in Luxembourgish
  2. (b)
    A quantitative and qualitative error analysis
  3. (c)
    An analysis on the usefulness of language exams as a tool for performance prediction on text generation tasks.
  4. (d)
    An accessible public leaderboard that can be used to recommend LLMs for given use cases [^2] The leaderboard can be found at https://applux-nnjdvksvbpgovzux7dtqts.streamlit.app/
2 Methodology
In this section, we give details on the methodology of this study and on our experiments, including the datasets we used, the LLMs we targeted, the metrics to compare the LLMs, and the prompt design. We also present and elaborate on the research questions we address in this paper.
We evaluate LLMs in two general disciplines: (1) language fluency and (2) language generation. The language fluency test consists of solving language exams that comprise sets of multiple-choice questions that relate to the general understanding of the language itself. In order to evaluate language generation, we use two tasks from the LuxGen benchmark Plum et al. (2025), specifically the headline generation and short description tasks. These tasks both consist of processing texts and generating appropriate summaries, either in form of a short description or a news headline.
Figure 1 shows the general testing pipeline. We first initiate the model with a unified prompt across every LLM (1). This initiation prompt consists of explaining the LLM’s role and the general task it needs to solve, as well as specifying the format of the inputs and outputs. We then pre-process the test samples to make them uniform and compatible with our prompts (2), and give them to the LLM in random order (3). For the language exams, we apply a post-processing step before comparing it to the ground truth (4). Post-processing generally consists of correcting minor typos in case the LLM’s output is different from each of the provided answers (however, we judge outputs as being invalid if they deviate significantly or are ambiguous and could be attributed to multiple answers). Finally, the LLM’s performance will be measured using task-specific metrics (see Section 2.4).
Figure 1: General Testing Pipeline
2.1 Research Questions
In this study, we aim to answer the following three research questions:
  1. RQ1: How well do LLMs perform in Luxembourgish language tests? To answer this question, we let LLMs solve language exams. We compare their performances in terms of number of correctly answered questions and determine a "winner".
  2. RQ2: In which cases do LLMs fail to understand Luxembourgish? For this question, we categorise the exam questions to determine which categories are more difficult for LLMs to solve.
  3. RQ3: Is there a correlation between performing well in a language exam and performing well in a language generation task? For the final question, we determine whether using language exams is an appropriate method to decide an LLM’s language capabilities and if the "winners" of the exams manage to perform well in generation tasks.
Category A1 A2 B1 B2 C1 C2 Total
Vocabulary 26 26 25 27 26 25 155
Grammar 26 26 26 33 28 26 165
Reading C. 26 26 25 26 26 26 155
Conversation C. 26 26 26 26 25 26 155
Total 104 104 102 112 105 103 630
Table 1: Statistics for language tests
2.2 Datasets
2.2.1 Language Tests
In order to evaluate the fluency of each model, we use official language exams for Luxembourgish provided to us by the Institut National des Langues Luxembourg, a language institute in Luxembourg [^3] https://www.inll.lu/en/. The difficulty level of each test follows the Common European Framework of Reference for Languages (CEFR) guidelines, and thusly ranges from A1 (beginner) to C2 (proficient) [^4] Further explanations can be found at https://www.efset.org/cefr/. The tests consist of mixtures of multiple-choice questions (MCQs) and open fill-in-the-blank type questions. However, for our experiments, we unify the exams by transforming them into the same type of questions. One of the authors who speaks the language at a native level converted the open questions into MCQs as this allows to use a single standardised prompt across all tests, and simplifies the error analysis. Questions fall into one of four broad categories: vocabulary, grammar, reading comprehension (RC), and listening comprehension. Note that the data for the latter category consists of transcripts rather than audio files. As such, we will refer to that category as conversation comprehension (CC) in this paper. Table 1 shows further details for the tests.
2.2.2 News Headline Generation
The headline generation task, being part of Plum et al.’s (2025) LuxGen benchmark for assessing Luxembourgish language models, consists of generating an appropriate title for a given news article. The articles were sourced from the Luxembourgish news agency RTL. The original test set comprises 13 852 samples, but for our purposes, we use a subset of 300 randomly selected samples due to the vast number of LLMs we are investigating.
2.2.3 Short Description Generation
The second LuxGen task we consider involves generating short descriptions of Luxembourgish Wikipedia articles. The descriptions featured on Wikidata serve as a ground truth for this task. The original test set consists of 2094 articles, but similarly to the headline generation task, we limit ourselves to a random selection of 300 samples.
Model Size A1 A2 B1 B2 C1 C2 Model Size A1 A2 B1 B2 C1 C2
Small Models ($\leq 15B$) Medium-sized Models ($>15B\&\leq 200B$)
Qwen 2 0.5 26.9 22.1 23.3 22.3 19.0 18.4 Mistral-Small 22 4.8 3.8 2.9 0.0 1.9 1.0
Llama 3.2 1 1.9 3.8 1.0 4.5 11.4 3.9 Gemma 2 27 21.2 6.7 3.9 1.8 2.9 2.9
Qwen 2 1.5 32.7 27.9 26.2 32.1 30.5 22.3 Aya-23 35 7.7 3.8 1.0 3.6 4.8 1.0
StableLM 2 1.6 0.0 14.4 0.0 17.0 9.5 15.5 Command-R 35 6.7 1.9 1.0 0.9 1.9 0.0
Gemma 2 2 39.4 32.7 30.1 0.0 36.2 28.2 Alfred 40 1.0 0.0 1.0 0.0 3.8 1.0
Llama 3.2 3 37.5 35.6 28.2 36.6 31.4 23.3 Mixtral 8x7 1.0 0.0 0.0 0.0 0.0 1.0
Phi 3 3.8 45.2 36.5 27.2 26.8 26.7 17.5 Llama 2 70 1.9 0.0 0.0 0.0 0.0 1.0
Phi 3.5 3.8 6.7 12.5 1.9 1.8 16.2 21.4 Llama 3 70 17.3 11.5 9.7 9.8 10.5 1.9
Qwen 2 7 48.1 36.5 34.0 35.7 36.2 33.0 Llama 3.1 70 21.2 12.5 33.0 10.7 11.4 1.9
Llama 2 7 7.7 10.6 8.7 13.4 23.8 20.4 DeepSeek-R1 70 17.3 13.4 9.7 7.1 8.6 7.8
WizardLM 2 7 24.0 30.8 27.2 19.6 35.2 23.3 Qwen 2 72 19.2 9.6 2.9 4.5 1.9 2.9
Mistral 7 35.6 35.6 35.0 31.3 31.4 26.2 Command-R+ 104 13.5 9.6 2.9 1.8 5.7 1.0
Aya-23 8 40.4 30.8 29.1 31.3 28.6 33.0 Mistral-Large 123 16.3 5.8 6.8 5.4 7.6 3.9
Llama 3.1 8 56.7 33.7 33.0 33.9 40.0 33.0 Mixtral 8x22 2.9 0.0 0.0 0.0 1.9 1.9
Llama 3 8 45.2 35.6 35.0 35.7 31.4 31.1 WizardLM 2 8x22 9.6 3.8 3.9 0.9 1.9 1.9
DeepSeek-R1 8 42.3 31.7 35.0 41.1 26.7 26.2 Large Models ($>200B$)
GLM 4 9 54.8 41.3 38.8 45.5 38.1 34.0 Llama 3.1 405 78.8 39.4 49.5 59.8 56.2 47.6
Gemma 2 9 58.7 47.1 39.8 0.0 37.1 33.0 DeepSeek-R1 671 92.3 83.7 74.8 76.8 72.4 82.5
Mistral-Nemo 12 52.9 36.5 30.1 34.8 28.6 21.4 ChatGPT 3.5 unk. 92.3 85.6 81.6 83.9 77.1 65.0
StableLM 2 12 27.9 7.7 26.2 29.5 25.7 24.3 Claude 3.5 Sonnet unk. 92.3 92.3 81.6 88.4 77.1 68.9
Llama 2 13 27.9 0.0 0.0 0.0 22.9 23.3 ChatGPT 4o unk. 92.3 85.6 79.6 84.8 74.3 60.2
Phi 3 14 45.2 38.5 39.8 39.3 35.2 24.3 ChatGPT 4o mini unk. 90.4 90.4 60.2 59.8 55.2 43.7
Gemini 2.0 Flash unk. 89.4 84.6 75.7 81.3 75.2 54.4
LeChat unk. 42.3 53.8 39.8 52.7 41.9 32.0
Table 2: Total results of language exams. Bold numbers indicate highest performance for each size category.
2.3 Models
We perform our experiments on a large number of LLMs from a wide array of language model families and model sizes. The main criteria for including a given LLM in our selection is their reported ability to (a) understand Luxembourgish, (b) understand a closely related Germanic language such as German or Dutch, or (c) understand a very large number of languages even if Luxembourgish is not explicitly included. Specifically, we include the following models: Alfred Hallström et al. (2023), Aya-23 Aryabumi et al. (2024), ChatGPT (OpenAI, 2023, OpenAI, 2024a, OpenAI, 2024b), Claude 3.5 Sonnet Anthropic (2024), Command-R Cohere For AI (2024), DeepSeek-R1 DeepSeek-AI (2025), Gemini 2.0 Flash Google , LeChat Mistral AI (2023a), Gemma 2 Team (2024), GLM 4 GLM et al. (2024), Llama 2 Touvron et al. (2023), Llama 3, Llama 3.1, Llama 3.2 Grattafiori et al. (2024), Mistral Mistral AI (2023b), Mixtral Mistral AI (2024), Phi 3, Phi 3.5 Abdin et al. (2024), Qwen 2 qwe (2024), StableLM 2 Bellagente et al. (2024), and WizardLM 2 Xu et al. (2024). Appendix B shows the full list of LLMs together with model sizes and reasons for inclusion.
Model sizes range from 1 billion parameters (Llama3.2:1B) to 671 billion (DeepSeek-R1:671B). We divide them into 3 size categories: small ($\leq 15B$ parameters), medium-sized ($>15B$& $\leq 200B$ parameters), and large ($>200B$ parameters). Note that although some model sizes are unknown to us as they are not communicated by the companies releasing them, they are estimated to be at least hundreds of billions of parameters. As such, we categorise the concerned LLMs as large.
2.4 Metrics
We measure the performance of LLMs using different metrics depending on the task. For the language exams, we use simple accuracy, i.e., the percentage of questions that an LLM answers correctly.
For language generation, we evaluate the outputs of LLMs in terms of grammar and spelling as well as similarity to the ground truth. We measure the performance on both of these tasks in an LLM-as-a-judge fashion using an LLM that is not part of the ones we are investigating as the judging LLM. We choose the judging LLM based on its ability to solve the task at hand, which we manually verify. The metrics used should meet the following requirements:
  • Grammar and Spelling metric
    1. 1.
      range between 00 and $1$ where 00 is incomprehensible and $1$ is perfect with neither spelling nor grammar mistakes.
    2. 2.
      punish non-Luxembourgish outputs, i.e., return 00 if the output is in a language other than Luxembourgish.
  • Ground Truth Similarity metric
    1. 1.
      range between 00 and $1$ where $1$ is identical to the ground truth and decreases the less similar the output is to the ground truth and the more it deviates from the given text.
    2. 2.
      punish descriptions that are deemed too long.
Considering that such LLM-generated metrics are essentially black boxes and non-deterministic due to the inherent random nature of LLMs, we repeat the generation a total of three times in order to reduce variations.
2.5 Prompt Design
We experimented with prompts of different lengths and degrees of complexity to find a prompt that yields the best for solving the language exams. See Appendix C for the full prompt. Regarding the prompts for the headline generation and description summary tasks, we reuse the prompts used by Plum et al. (2025) for every model we investigate. The prompts used for evaluating the LuxGen tasks that uses the metric requirements described in Section 2.4 can be found in Appendix D.
3 Results
3.1 RQ1: How well do LLMs perform in Luxembourgish language tests?
Table 2 shows the overall results of the LLMs we tested. See Appendix E for a breakdown of results across the four categories. For a more convenient way of exploring the results, we also release an interactive web app with the results 555which can be found at https://applux-nnjdvksvbpgovzux7dtqts.streamlit.app/.
We observe a clear trend where the large models typically perform well above average. Claude 3.5 Sonnet achieves the highest scores across every test except for the C2 test for which DeepSeek-R1:671B performs best. Among the small models, many models show promise to be adequate for further fine-tuning on Luxembourgish data despite their small size as they either approach or slightly surpass the 50% mark on the easier tests. Gemma2:9B performs best on the easier tests, but we deem GLM4 to be the best small model overall as it gets close to Gemma2’s scores for the easier tests and tends to perform best on the harder tests.
There is also a trend appearing of small models outperforming medium-sized models, with the latter performing worse than guessing at random. Llama3.1:70B is the overall winner among the medium-sized models, however, it performs significantly worse than GLM4, and indeed most of the small models.
Furthermore, as expected, performances drop for higher CEFR levels as the tests become increasingly more difficult, with one notable exception. Curiously, most models achieve a higher performance on the B2 test than on the B1 test. Specifically, they perform surprisingly well on the Reading Comprehension questions, however, it is unclear as to why those questions appear to be easier for LLMs to solve.
3.2 RQ2: In which cases do LLMs fail to understand Luxembourgish?
In order to answer this question, we analyse trends and patterns relating to the four categories mentioned in Section 2.2.1. Table 3 contains the performances of the studied language models per category. From the results, we note that overall, LLMs perform almost consistently best in vocabulary questions while grammar appears to be the hardest category. Furthermore, although the Reading Comprehension and Conversation Comprehension questions are nominally different, there is no statistical difference between the performances of the LLMs in either category (according to a paired t-test). However, when taking into account the size categories of the LLMs, large models perform significantly better in the reading comprehension category than conversation comprehension questions.
In addition to the coarse-grained ones, we created fine-grained categories for the questions in the language tests. Table 4 shows the categories as well as the number of questions that fit into the categories. We present the results of the best-performing models in Table 5. Due to space issues, we put the entire table in Appendix F. The categorisation was done in a supervised fashion using Claude 3.7 Sonnet and a separate manually annotated language exam. We provide the descriptions for each category as well as the prompt for annotation in Appendices G and H. We manually reviewed the automatically annotated questions to ensure correctness.
model Size Vocab. Grammar RC CC
Small Models ($\leq 15B$)
Qwen 2 0.5 26.5 21.8 19.4 21.9
Llama 3.2 1 10.3 4.8 1.3 1.3
Qwen 2 1.5 34.8 32.1 23.9 25.2
StableLM 2 1.6 14.8 13.9 5.2 4.5
Gemma 2 2 29.0 24.8 27.7 29.7
Llama 3.2 3 33.5 27.9 38.7 32.3
Phi 3 3.8 27.7 26.1 34.8 34.8
Phi 3.5 3.8 14.8 4.8 11.6 8.4
Qwen 2 7 34.2 32.7 45.2 38.7
Llama 2 7 12.3 6.7 14.8 26.5
WizardLM 2 7 20.6 26.7 29.0 28.4
Mistral 7 30.3 27.3 36.8 38.1
Aya-23 8 30.3 28.5 39.4 35.5
Llama 3.1 8 32.9 34.5 50.3 39.4
Llama 3 8 34.2 36.4 38.7 38.1
DeepSeek-R1 8 38.1 30.9 32.9 37.4
GLM 4 9 50.3 24.8 51.6 48.4
Gemma 2 9 43.9 24.8 41.3 34.2
Mistral-Nemo 12 37.4 27.9 40.0 35.5
StableLM 2 12 27.1 14.5 31.0 27.7
Llama 2 13 18.1 14.5 12.9 14.8
Phi 3 14 36.1 37.6 38.7 37.4
Medium-sized Models ($>15B\&\leq 200B$)
Mistral-Small 22 7.7 1.2 0.0 0.6
Gemma 2 27 12.3 8.5 2.6 2.6
Aya-23 35 9.0 1.8 1.3 3.2
Command-R 35 5.2 0.6 0.6 1.9
Alfred 40 1.3 1.8 0.0 1.3
Mixtral 8x7 0.6 0.6 0.0 0.0
Llama 2 70 0.0 1.2 0.0 0.6
Llama 3 70 22.6 12.7 1.9 3.2
Llama 3.1 70 26.5 14.5 2.6 5.2
DeepSeek-R1 70 20.6 18.2 0.6 2.6
Qwen 2 72 12.3 7.3 2.6 5.2
Command-R+ 104 14.8 3.0 2.6 2.6
Mistral-Large 123 21.3 6.7 1.3 1.3
Mixtral 8x22 1.9 1.8 0.0 0.6
WizardLM 2 8x22 7.7 4.8 0.6 1.3
Large Models ($>200B$)
Llama 3.1 405 64.5 43.6 62.6 53.5
DeepSeek-R1 671 93.5 73.9 84.5 76.1
ChatGPT 3.5 unk. 92.9 72.1 87.1 80.0
Claude 3.5 Sonnet unk. 89.0 81.8 88.4 81.9
ChatGPT 4o unk. 89.7 71.5 84.5 80.6
ChatGPT 4o mini unk. 78.1 55.2 72.9 65.8
Gemini 2.0 Flash unk. 83.2 67.3 83.2 79.4
LeChat unk. 61.3 34.5 40.0 45.8
Table 3: Total results of language exams. Bold numbers indicate highest performance for each size category.
Category name Number
V_idiom 21
V_noun 84
V_verb 44
V_adjective 23
V_native 6
G_gender 37
G_pronouns 20
G_verbs 55
G_adjective 9
G_tense 50
G_past 21
G_present 13
G_subjunctive 14
G_modal 7
G_preposition 54
G_cases 39
RC_TF 152
RC_Simple 72
RC_Complex 86
Table 4: Statistics for fine-grained categories for exam questions.
It is to note that for this categorisation, we do not differentiate between reading comprehension and conversation comprehension, and thusly group them together.
For vocabulary-related questions, there is a clear pattern showing that LLMs perform worse in categories that require "language-exclusive" knowledge and cannot be easily inferred from other languages, i.e., idioms and native words. Regarding grammar, the hardest categories for the large models are adjectives, past tense, and cases. The main challenge appears to be related to the fact that the possible answers for many of those questions are almost identical. For example, one question from the adjectives category is:
Op ___ Weiere kann een de Wanter Schlittschong fueren
with the possible answers being:
zougefruer, zougefruerene, zougefruerenen, zougefréieren
For the reading comprehension questions, we split them into three categories: true/false, simple, and complex. Simple questions do not require an understanding of the entire given text excerpt, and the correct answer for those questions can be found either explicitly written in the text or by understanding a small portion of the text. True/false and complex questions require a model to fully understand the given text excerpt, the main difference being that True/False questions are binary while complex ones have three or four possible answers. As such, the latter questions are considered harder. This assumption is indeed confirmed by the results as the large models seem to perform best for simple questions.
It is to note, that these observations largely hold true for the best-performing small models, but performances tend to fluctuate for worse-performing models as well as for medium-sized models.
model Llama 3.1 405B DeepSeek-R1 671B ChatGPT 3.5 Claude 3.5 ChatGPT 4o ChatGPT 4o m Gemini 2.0 LeChat
V_Idioms 42.9 81.0 71.4 42.9 52.4 38.1 33.3 38.1
V_Noun 67.9 94.0 91.7 85.7 86.9 79.8 81.0 65.5
V_Verb 68.2 93.2 90.9 88.6 90.9 68.2 84.1 61.4
V_Adjective 47.8 91.3 100.0 100.0 95.7 87.0 87.0 47.8
V_Native 33.3 100.0 50.0 50.0 66.7 16.7 16.7 50.0
G_Gender 37.8 70.3 73.0 70.3 78.4 56.8 70.3 24.3
G_Pronouns 40.0 65.0 60.0 75.0 60.0 65.0 70.0 50.0
G_Verb 41.8 69.1 74.5 89.1 74.5 52.7 72.7 34.5
G_Adjective 22.2 55.6 44.4 66.7 55.6 22.2 66.7 44.4
G_Tense 38.0 70.0 76.0 90.0 76.0 54.0 76.0 34.0
G_Past 33.3 57.1 71.4 81.0 66.7 47.6 66.7 28.6
G_Present 53.8 84.6 92.3 100.0 92.3 76.9 84.6 38.5
G_Subjunctive 28.6 71.4 71.4 92.9 78.6 42.9 71.4 35.7
G_Modal 14.3 100.0 71.4 85.7 85.7 71.4 85.7 42.9
G_Prepositions 55.6 81.5 75.9 85.2 74.1 64.8 68.5 35.2
G_Cases 33.3 64.1 69.2 69.2 64.1 51.3 53.8 35.9
RC_TF 55.9 79.6 85.5 88.2 83.6 61.8 80.9 39.5
RC_Simple 63.6 86.4 88.6 86.4 86.4 86.4 86.4 40.9
RC_Complex 59.1 84.8 83.3 86.4 83.3 81.8 84.8 47.0
Table 5: Performance of large models on fine-grained categories. V: Vocabulary ; G: Grammar ; RC: Reading Comprehension ; TF: True/False
3.3 RQ3: Are there correlations between performing well in a language exam and performing well in a language generation task?
Task Small Medium Large All
Description - Grammar 0.664 0.765 0.782 0.385
Description - Adequacy 0.645 0.724 0.622 0.267
Headline - Grammar 0.486 0.552 0.093 0.352
Headline - Adequacy 0.644 0.598 0.269 0.284
Table 6: Pearson Correlation Coefficient between the overall exam scores and LuxGen task performances
For this research question, we first average out the performances of every LLM on all six language exams, resulting in their overall exam performance score. We then let the LLMs solve both of the LuxGen tasks, and evaluate the outputs using an LLM-as-a-judge. Our LLM of choice is Claude 3.7 Sonnet Anthropic (2025). After manually verifying its capabilities to solve the tasks at hand, we deemed it an adequate LLM to evaluate the experiments as it solved both tasks well and with minimal spelling mistakes. For both tasks and both metrics we employ (see Section 2.4), we calculate mean performance scores for each LLM. Figure 2 shows the performances of all LLMs on the LuxGen tasks ($y$-axis) plotted against their respective overall exam performances ($x$-axis). See Appendix I for detailed results. We observe the formation of well-defined clusters, each corresponding to a different LLM size category which remains relatively consistent across each task and metric. Generally, the LuxGen performances of the small and large LLMs appear to strongly positively correlate with their exam performances. However, the medium-sized LLMs seem to not follow the same trend. Despite their generally low exam scores, many of them perform better on LuxGen tasks than other LLMs that reached high exam scores. However, when considering each cluster individually, we largely observe strong positive correlations that indeed show higher LuxGen performances for higher exam scores.
To confirm our observations, we calculate Pearson’s Correlation Coefficients (PCCs) for each set of performances (see Table 6). Across the entire set of LLMs, the PCCs are positive and weak to moderate, while within clusters, the PCCs typically show that performances are highly positively correlated. One notable exception is the large category of LLMs for the News Headline generation task, where there appears to be either a weak positive correlation (for adequacy) or no correlation at all (for grammar and spelling).
(a) Description Task - Grammar and Spelling
(b) Description Task - Adequacy
(c) Headlines Task - Grammar and Spelling
(d) Headline Task - Adequacy
Figure 2: Performance on Language Exams vs Performance on the LuxGen Tasks
4 Discussion
Generally, we conclude that language exams can indeed be a useful tool to assess the generation capabilities of LLMs. For the NLP tasks we tested, we can to a degree predict if an LLM performs well given how well it scored on the exams. However, there are some caveats to keep in mind:
First, for a considerable number of LLMs, these exams are not an adequate method. LLMs that have between 15 billion and 200 billion parameters scored consistently low on the exams, yet they performed well in the LuxGen tasks. Secondly, it is important to consider for what type of task an LLM is considered. Language exams are used to evaluate the general proficiency in a language across multiple disciplines. As such, the overall exam score can be dragged down by irrelevant categories of questions, e.g., an LLM used to detect hate speech in an online forum does not necessarily need to have a perfect grasp on obscure grammar rules. In such a case, it is preferable to adjust the exam to only include certain question categories. Thirdly, for this study, we did not do any additional fine-tuning on Luxembourgish data which might have altered the performances of the tested LLMs to different degrees. It is possible that certain LLMs adapt better when fine-tuned on a Luxembourgish instruction tuning dataset.
5 Related Works
5.1 Testing LLMs on Exams
Many recent studies have used exams to test the knowledge of LLMs in a variety of domains. These domains predominantly cover the medical domain (Gotta et al., 2024, Stribling et al., 2024, Şahin et al., 2025, Abbas et al., 2024,Zong et al., 2024), but also computer science and programming (Dinh et al., 2024, Ellis et al., 2024, Li et al., 2025,VarastehNezhad et al., 2024), and the legal domain (Katz et al., 2024, Martínez, 2024). The general consensus is that LLMs perform well in these exams, often surpassing human candidates. However, most of these studies cover exams written in English, with only a few covering exams from Non-English speaking countries. Koto (2024) released IndoCareer, an exam benchmark covering various topics in Indonesian. They test it on a large number of LLMs with mixed results. Koto et al. (2023) also conducted a similar study using Indonesian school exams, where they found that most LLMs managed to only pass primary school exams. Jassem et al. (2025) presented LLMzSzŁ, a benchmark of Polish national exams covering various subjects. They test numerous LLMs in both English and Polish, showing that most LLMs perform better than randomly guessing. Locatelli et al. (2024) tested LLMs on a standardised entry exam for Brazilian universities covering different high school subjects. They found that all LLMs performed very well on the exams, however, the number of tested LLMs was very low compared to our study. Dao et al. (2023) compared two LLMs using Vietnamese high school physics exams, showing that they perform worse than students.
5.2 Testing LLMs on Language Exams
While there are numerous studies on the performance of LLMs on general NLP tasks, they are far less commonly focused specifically on language exams. Darg̀is et al. (2024) systematically test numerous LLMs on standardised high school language tests in Latvian and manually evaluate their outputs with a human expert. Mayor-Rocher et al. (2024) tested several LLMs on TELEIA, a benchmark of exams for Spanish as a second language, as well as a Spanish test for foreign university students. Similarly, Mercorio et al. (2024) test numerous LLMs on the novel Italian language exam benchmark INVALSI, finding that large models consistently outperform smaller ones, and that the performance degrades on harder tests.
However, compared to related studies, we do not only assess LLM performance on language exams alone, but we also explore their usefulness to predict their performance in general NLP tasks.
6 Conclusion
In this study, we investigated the performance of a large number of recent LLMs on official language exams, and assessed whether or not they serve as a useful tool to determine the capabilities of language models. We found that many models struggle with even the easier exams, with only the largest models scoring well in every exam. We also delved into dividing the exam questions into fine-grained categories to assess which types of questions are harder to solve for LLMs, finding that questions requiring language-exclusive knowledge as well as questions with similar answers pose problems to many models. Finally, we determined that using such exams can indeed be used as an indicator for text generation and understanding capabilities of LLMs, with some caveats to consider.
7 Acknowledgments
This work is supported by the LLMs4EU project, funded by the European Union through the Digital Europe Programme (DIGITAL) under the grant agreement 101198470, and by the BESSER project through the Luxembourg National Research Fund (FNR) PEARL programme under the grant agreement 16544475.
We would like to thank the Institut National des Langues Luxembourg for providing the Luxembourgish exams for our study.
A Limitations
While we are confident about the findings and conclusions of our studies, there are still certain limitations to consider. First, while we aimed to include a comprehensive list of current LLMs in our experiments, it is a nearly impossible task considering the frequent release of new models. We even had to scrap parts of our experiments that were done using Gemini 1.5 Flash as it became inaccessible after the release of the 2.0 model. We are also fully aware that many new model versions have been released that we did not consider for this study. We however remain confident that our selection of LLMs gives a good overview of available models. Another limitation relates to the relatively small number of generation tasks we considered for RQ3. As they are essentially both summarisation tasks, it is possible that our conclusions do not apply to other types of NLP tasks. This could be a useful research direction for future work. Finally, it is possible that our conclusions are not generalisable to other languages since we are limiting ourselves to Luxembourgish. Again, studying if we can apply our methodology to other low-resource languages would be a worthwhile research direction that can benefit other language communities.
B Full List of LLMs and Reasons for Inclusion
Table 7 contains the full list of LLMs we include in this study.
Model Size Reason for Inclusion
Small Models $(\leq 15B)$
Qwen 2 0.5 Supports German and Dutch
Llama 3.2 1 Supports German
Qwen 2 1.5 Supports German and Dutch
StableLM 2 1.6 Supports German and Dutch
Gemma 2 2 Was trained on multilingual data
Llama 3.2 3 Supports German
Phi 3 3.8 Tested on multilingual NLP tasks
Phi 3.5 3.8 Tested on multilingual NLP tasks
Qwen 2 7 Supports German and Dutch
Llama 2 7 Was trained on multilingual data / not explicitly multilingual
WizardLM 2 7 claims to be multilingual
Mistral 7 European model / not explicitly multilingual
Aya-23 8 Supports German and Dutch
Llama 3.1 8 Supports German
Llama 3 8 Performed well on multilingual NLP task including German
DeepSeek-R1 8 smaller version of massively multilingual model
GLM 4 9 Supports German
Gemma 2 9 Training on multilingual data
Mistral-Nemo 12 Supports German
StableLM 2 12 Supports German and Dutch
Llama 2 13 Was trained on multilingual data
Phi 3 14 Tested on multilingual NLP tasks
Medium-sized Models $(>15B\&\leq 200B)$
Mistral Small 22 Supports German and Dutch
Gemma 2 27 Training on multilingual data
Aya-23 35 Supports German and Dutch
Command-R 35 Supports German
Alfred 40 Supports German
Mixtral 8x7 Supports German
Llama 2 70 Was trained on multilingual data
Llama 3 70 Performed well on multilingual NLP task including German
Llama 3.1 70 Supports German
DeepSeek-R1 70 smaller version of massively multilingual model
Qwen 2 72 Supports German and Dutch
Command-R+ 104 Supports German
Mistral-Large 123 Supports German
Mixtral 8x22 Supports German
WizardLM 2 8x22 claims to be multilingual
Large Models $(>200B)$
Llama 3.1 405 Supports German
DeepSeek-R1 671 Massively Multilingual Model
ChatGPT 3.5 175 Massively Multilingual Model
Claude 3.5 Sonnet unk Massively Multilingual Model
ChatGPT 4o unk Massively Multilingual Model
ChatGPT 4o mini unk Massively Multilingual Model
Gemini 2.0 Flash unk Massively Multilingual Model
LeChat unk Massively Multilingual Model
Table 7: Full list of LLMs and our reasoning for including them in this study
C Language Exam Prompt
We use the following prompt for testing LLMs on language exams:
I will give you a language test for
Luxembourgish. For each part, you will
get a TEXT with one part missing marked
by [BLANK] and a list of possible ANSWERS
where every possible option is separated
with a comma. You have to choose the
option that best fits as replacement for
the given BLANK. As output, simply write
the chosen option, nothing else.
INPUT: [the question and possible answers]
OUTPUT:
where the INPUTs for the exams are formatted as follows:
TEXT: Den Direkter ass houfereg [BLANK]
eis Leeschtung.Ψ
ANSWERS: [iwwert,fir,mat,op]
D Evaluation Metric Prompts for LuxGen
For each task and metric, we designed different prompts for a total of four prompts:
  • Grammar and Spelling Metric for Small Description Task:
    I will give you a table as a TSV file. This
    table contains Luxembourgish Wikipedia
    articles in the first column, the second
    column contains ground truth short
    descriptions for the articles. The rest of
    the columns contain short descriptions
    generated by various language models. The
    generated descriptions are supposed to be
    written in correct Luxembourgish. Your
    task is to judge the generated descriptions
    in terms of grammatical correctness and
    spelling. Keep in mind that the descriptions
    do not need to be complete sentences.
    For each language model and sample, return
    a score ranging from 0 to 1 with two
    decimals behind the comma. For the score,
    1 means that the description is perfectly
    written without any grammatical or spelling
    mistakes. The more mistakes there are, the
    lower the score should be. A 0 means that
    the headline is full of mistakes and made
    up words. Also return 0 if the description
    is not written in Luxembourgish at all.
    Return the average scores in a table.
  • Adequacy Metric for Small Description Task:
    I will give you a table as a TSV file. This
    table contains Luxembourgish Wikipedia
    articles in the first column, the second
    column contains ground truth short
    descriptions for the articles. The rest of
    the columns contain short descriptions
    generated by various language models. The
    generated descriptions are supposed to be
    written in correct Luxembourgish. Your
    task is to judge the generated
    descriptions in terms of the content
    being an adequate description of the
    article. Keep in mind that the
    descriptions do not need to be complete
    sentences.
    For each language model and sample, return
    a score ranging from 0 to 1 with two
    decimals behind the comma. For the score,
    1 means that the description is identical
    to the ground truth, and should be lower
    the more it deviates from the meaning of
    the article. The score should also be
    lowered for descriptions that are too long.
    A 0 means that the description is
    completely irrelevant to the article.
    Return the average scores in a table.
  • Grammar and Spelling Metric for News Headline Task:
    You are a news editor for a Luxembourgish
    news paper. I will give you a table as a
    TSV file. This table contains news
    articles in the first column, the second
    column contains ground truth headlines
    for the articles. The rest of the columns
    will contain headlines that were produced
    by various language models. Your task is
    to judge the produced headlines in terms
    of grammatical correctness and spelling.
    For each language model and sample, return
    a score ranging from 0 to 1 with two
    decimals behind the comma. For the score,
    1 means that the headline is perfectly
    written without any grammatical or
    spelling mistakes. The more mistakes there
    are, the lower the score should be. A 0
    means that the headline is full of
    mistakes and made up words. Also return
    0 if the description is not written in
    Luxembourgish at all.
    Return the average scores in a table.
  • Adequacy Metric for News Headline Task:
    You are a news editor for a Luxembourgish
    news paper. I will give you a table as a
    TSV file. This table contains news
    articles in the first column, the second
    column contains ground truth headlines
    for the articles. The rest of the columns
    will contain headlines that were produced
    by various language models. Your task is
    to judge the produced headlines in terms
    of summary of the content of the article.
    For each language model and sample,
    return a score ranging from 0 to 1 with
    two decimals behind the comma. For the
    score, a 1 means that headline is
    identical to the ground truth, and should
    decrease the more its meaning deviates
    from the content in the news article. The
    score should also be lowered for headlines
    that are too long. A 0 means that the
    headline is completely irrelevant to the
    article.
    Return the average scores in a table.
E Coarse-grained Exam Results
Table 8 shows the full breakdown of our results on language tests for each language level.
A1 A2 B1 B2 C1 C2
Model Size Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC Voc. Gra. RC CC
Small Models $(\leq 15B)$
Qwen 2 0.5 38.5 26.9 23.1 19.2 34.6 11.5 19.2 23.1 32.0 15.4 23.1 23.1 25.9 15.2 19.2 30.8 19.2 28.6 15.4 20.0 8.0 34.6 11.5 19.2
Llama 3.2 1 0.0 0.0 3.8 3.8 11.5 3.8 0.0 0.0 4.0 0.0 0.0 0.0 11.1 3.0 0.0 3.8 30.8 14.3 0.0 0.0 4.0 7.7 3.8 0.0
Qwen 2 1.5 34.6 38.5 42.3 15.4 38.5 26.9 23.1 23.1 32.0 30.8 19.2 23.1 33.3 36.4 23.1 34.6 50.0 39.3 11.5 32.0 20.0 19.2 23.1 26.9
StableLM 2 1.6 0.0 0.0 0.0 0.0 19.2 23.1 0.0 15.4 0.0 0.0 0.0 0.0 25.9 24.2 11.5 3.8 19.2 7.1 0.0 12.0 24.0 26.9 0.0 11.5
Gemma 2 2 38.5 46.2 38.5 34.6 38.5 23.1 34.6 34.6 40.0 19.2 30.8 30.8 0.0 0.0 0.0 0.0 19.2 42.9 42.3 40.0 40.0 23.1 23.1 26.9
Llama 3.2 3 34.6 34.6 46.2 34.6 38.5 23.1 26.9 53.8 40.0 26.9 23.1 23.1 37.0 27.3 42.3 42.3 38.5 25.0 23.1 40.0 12.0 30.8 30.8 19.2
Phi 3 3.8 26.9 46.2 57.7 50.0 34.6 23.1 38.5 50.0 24.0 23.1 23.1 38.5 25.9 27.3 15.4 38.5 30.8 28.6 23.1 24.0 24.0 7.7 19.2 19.2
Phi 3.5 3.8 7.7 0.0 19.2 0.0 11.5 7.7 15.4 15.4 0.0 3.8 3.8 0.0 0.0 3.0 0.0 3.8 30.8 3.6 19.2 12.0 40.0 11.5 15.4 19.2
Qwen 2 7 42.3 26.9 69.2 53.8 23.1 38.5 38.5 46.2 36.0 23.1 15.4 61.5 29.6 33.3 38.5 42.3 38.5 32.1 38.5 36.0 36.0 42.3 15.4 38.5
Llama 2 7 7.7 0.0 7.7 15.4 0.0 0.0 19.2 23.1 8.0 0.0 23.1 3.8 7.4 6.1 7.7 34.6 26.9 17.9 23.1 28.0 24.0 15.4 23.1 19.2
WizardLM 2 7 7.7 15.4 38.5 34.6 30.8 26.9 30.8 34.6 24.0 34.6 19.2 30.8 7.4 18.2 26.9 26.9 38.5 32.1 26.9 44.0 16.0 34.6 15.4 26.9
Mistral 7 30.8 26.9 46.2 38.5 23.1 30.8 42.3 46.2 44.0 34.6 26.9 34.6 29.6 24.2 26.9 46.2 34.6 21.4 30.8 40.0 20.0 26.9 30.8 26.9
Aya-23 8 38.5 38.5 46.2 38.5 30.8 19.2 30.8 42.3 32.0 19.2 26.9 38.5 25.9 30.3 26.9 42.3 26.9 28.6 26.9 32.0 28.0 34.6 38.5 30.8
Llama 3.1 8 50.0 61.5 57.7 57.7 30.8 30.8 26.9 46.2 16.0 38.5 19.2 57.7 33.3 24.2 53.8 26.9 38.5 28.6 50.0 44.0 28.0 26.9 30.8 46.2
Llama 3 8 42.3 42.3 50.0 46.2 30.8 34.6 26.9 50.0 40.0 30.8 30.8 38.5 25.9 36.4 34.6 46.2 42.3 39.3 15.4 28.0 24.0 34.6 42.3 23.1
DeepSeek-R1 8 53.8 50.0 34.6 30.8 42.3 23.1 30.8 30.8 36.0 26.9 34.6 42.3 44.4 42.4 42.3 34.6 30.8 14.3 23.1 40.0 20.0 26.9 23.1 34.6
GLM 4 9 57.7 42.3 69.2 50.0 34.6 19.2 42.3 69.2 60.0 15.4 26.9 53.8 59.3 27.3 50.0 50.0 53.8 21.4 34.6 44.0 36.0 23.1 26.9 50.0
Gemma 2 9 76.9 50.0 61.5 46.2 61.5 38.5 30.8 57.7 52.0 26.9 34.6 46.2 0.0 0.0 0.0 0.0 38.5 17.9 46.2 48.0 36.0 23.1 42.3 30.8
Mistral-Nemo 12 61.5 50.0 53.8 46.2 34.6 26.9 38.5 46.2 44.0 15.4 23.1 38.5 33.3 24.2 38.5 46.2 30.8 25.0 26.9 32.0 20.0 26.9 26.9 11.5
StableLM 2 12 42.3 15.4 30.8 23.1 0.0 0.0 7.7 23.1 44.0 11.5 19.2 30.8 22.2 18.2 50.0 30.8 34.6 17.9 19.2 32.0 20.0 23.1 30.8 23.1
Llama 2 13 34.6 38.5 15.4 23.1 19.2 11.5 15.4 11.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 30.8 21.4 19.2 20.0 24.0 19.2 26.9 23.1
Phi 3 14 42.3 42.3 50.0 46.2 38.5 42.3 30.8 42.3 44.0 38.5 30.8 46.2 48.1 33.3 34.6 42.3 30.8 39.3 34.6 36.0 12.0 30.8 19.2 34.6
Medium-sized Models $(>15B\&\leq 200B)$
Mistral Small 22 15.4 0.0 0.0 3.8 11.5 3.8 0.0 0.0 8.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 7.7 0.0 0.0 0.0 4.0 0.0 0.0 0.0
Gemma 2 27 42.3 23.1 3.8 15.4 15.4 11.5 0.0 0.0 0.0 7.7 0.0 7.7 3.7 3.0 0.0 0.0 7.7 3.6 0.0 0.0 4.0 3.8 3.8 0.0
Aya-23 35 15.4 7.7 0.0 7.7 7.7 0.0 0.0 7.7 4.0 0.0 0.0 0.0 7.4 3.0 0.0 3.8 15.4 0.0 0.0 4.0 4.0 0.0 0.0 0.0
Command-R 35 19.2 3.8 0.0 3.8 3.8 0.0 0.0 3.8 4.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 3.8 0.0 0.0 4.0 0.0 0.0 0.0 0.0
Alfred 40 0.0 0.0 0.0 3.8 0.0 0.0 0.0 0.0 0.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 7.7 3.6 0.0 4.0 0.0 3.8 0.0 0.0
Mixtral 56 3.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 0.0 0.0
Llama 2 70 0.0 3.8 0.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 0.0 0.0
Llama 3 70 38.5 23.1 0.0 7.7 30.8 7.7 0.0 7.7 24.0 11.5 0.0 3.8 14.8 12.1 0.0 11.5 26.9 14.3 0.0 0.0 0.0 7.7 0.0 0.0
Llama 3.1 70 46.2 19.2 0.0 19.2 19.2 19.2 0.0 11.5 36.0 19.2 0.0 3.8 18.5 15.2 0.0 7.7 34.6 10.7 0.0 0.0 4.0 3.8 0.0 0.0
DeepSeek-R1 70 34.6 26.9 0.0 7.7 26.9 26.9 0.0 0.0 24.0 15.4 0.0 0.0 11.1 9.1 0.0 7.7 19.2 14.3 0.0 0.0 8.0 19.2 3.8 0.0
Qwen 2 72 34.6 15.4 3.8 23.1 19.2 7.7 0.0 11.5 8.0 3.8 0.0 0.0 3.7 9.1 0.0 3.8 7.7 0.0 0.0 0.0 0.0 7.7 0.0 3.8
Command-R+ 104 30.8 3.8 3.8 15.4 26.9 3.8 0.0 7.7 4.0 3.8 0.0 3.8 3.7 3.0 0.0 0.0 19.2 3.6 0.0 0.0 4.0 0.0 0.0 0.0
Mistral-Large 123 38.5 19.2 0.0 7.7 15.4 3.8 0.0 3.8 20.0 7.7 0.0 0.0 18.5 0.0 3.8 0.0 26.9 3.6 0.0 0.0 8.0 7.7 0.0 0.0
Mixtral 176 7.7 0.0 0.0 3.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.8 3.6 0.0 0.0 0.0 7.7 0.0 0.0
WizardLM 2 176 19.2 11.5 0.0 7.7 7.7 3.8 0.0 3.8 12.0 3.8 0.0 0.0 3.7 0.0 0.0 0.0 3.8 3.6 0.0 0.0 0.0 7.7 0.0 0.0
Large Models $(>200B)$
Llama 3.1 405 92.3 57.7 80.8 84.6 26.9 38.5 34.6 57.7 76.0 46.2 26.9 50.0 70.4 45.5 65.4 61.5 76.9 39.3 57.7 52.0 44.0 34.6 61.5 50.0
DeepSeek-R1 671 96.2 84.6 96.2 92.3 100 65.4 76.9 92.3 92.0 76.9 42.3 88.5 92.6 57.6 84.6 76.9 80.8 64.3 73.1 72.0 100 100 73.1 57.7
ChatGPT 3.5 unk 100 76.9 100 92.3 100 84.6 69.2 88.5 100 92.3 38.5 96.2 92.6 75.8 80.8 88.5 88.5 60.7 76.9 84.0 76.0 42.3 76.9 65.4
Claude 3.5 S unk 100 80.8 100 88.5 96.2 88.5 88.5 96.2 100 92.3 38.5 96.2 88.9 84.8 84.6 96.2 84.6 71.4 80.8 72.0 64.0 73.1 69.2 69.2
ChatGPT 4o unk 96.2 84.6 100 88.5 100 73.1 80.8 88.5 100 84.6 38.5 96.2 96.3 78.8 80.8 84.6 80.8 57.1 73.1 88.0 64.0 50.0 65.4 61.5
ChatGPT 4o m unk 96.2 73.1 100 92.3 100 84.6 80.8 96.2 72.0 46.2 42.3 80.8 77.8 54.5 50.0 57.7 73.1 39.3 57.7 52.0 48.0 34.6 42.3 50.0
Gemini 2.0 F unk 88.5 76.9 96.2 96.2 96.2 65.4 80.8 96.2 92.0 84.6 42.3 84.6 92.6 66.7 80.8 88.5 84.6 60.7 80.8 76.0 44.0 50.0 61.5 61.5
LeChat unk 80.8 26.9 3.8 57.7 73.1 50.0 38.5 53.8 48.0 26.9 34.6 50.0 74.1 33.3 53.8 53.8 53.8 42.9 26.9 44.0 36.0 26.9 34.6 30.8
Table 8: Detailed results on experiments with coarse-grained categories
F Fine-grained Exam Results
Table 9 shows results of our experiments on fine-grained categories for all LLMs.
Model Size V_Idioms V_Noun V_Verb V_Adj. V_Native G_Gender G_Pron. G_Verb G_Adj. G_Tense G_Past G_Present G_Subj. G_Modal G_Prep. G_Cases RC_TF RC_S. RC_C.
Small Models $(\leq 15B)$
Qwen 2 0.5 9.5 26.2 22.7 34.8 16.7 16.2 20.0 21.8 0.0 24.0 23.8 30.8 21.4 14.3 25.9 20.5 41.4 0.0 1.5
Llama 3.2 1 19.0 15.5 6.8 0.0 0.0 0.0 0.0 3.6 0.0 4.0 9.5 0.0 0.0 0.0 9.3 0.0 2.0 0.0 0.0
Qwen 2 1.5 33.3 39.3 27.3 30.4 16.7 37.8 55.0 20.0 22.2 20.0 28.6 23.1 0.0 0.0 40.7 35.9 38.2 9.1 13.6
StableLM 2 1.6 23.8 16.7 18.2 4.3 16.7 8.1 5.0 12.7 11.1 14.0 9.5 15.4 14.3 14.3 18.5 10.3 4.6 0.0 4.5
Gemma 2 2 28.6 31.0 20.5 43.5 33.3 29.7 25.0 23.6 22.2 24.0 28.6 7.7 35.7 0.0 24.1 20.5 40.1 18.2 19.7
Llama 3.2 3 28.6 35.7 29.5 34.8 33.3 18.9 15.0 27.3 44.4 24.0 19.0 38.5 21.4 14.3 35.2 20.5 39.5 22.7 30.3
Phi 3 3.8 23.8 23.8 36.4 21.7 33.3 32.4 35.0 23.6 11.1 26.0 19.0 30.8 21.4 42.9 24.1 25.6 38.2 29.5 33.3
Phi 3.5 3.8 38.1 19.0 11.4 8.7 33.3 2.7 5.0 5.5 11.1 4.0 4.8 0.0 7.1 0.0 3.7 5.1 9.2 13.6 10.6
Qwen 2 7 33.3 39.3 31.8 21.7 50.0 40.5 45.0 25.5 33.3 28.0 23.8 30.8 21.4 42.9 27.8 48.7 48.0 40.9 37.9
Llama 2 7 19.0 16.7 6.8 8.7 33.3 0.0 0.0 7.3 11.1 8.0 4.8 0.0 21.4 0.0 9.3 0.0 31.6 9.1 7.6
WizardLM 2 7 23.8 21.4 25.0 8.7 33.3 32.4 15.0 25.5 33.3 24.0 33.3 15.4 35.7 28.6 25.9 28.2 33.6 20.5 25.8
Mistral 7 28.6 36.9 25.0 17.4 16.7 29.7 35.0 29.1 22.2 26.0 38.1 15.4 14.3 28.6 20.4 33.3 40.8 25.0 37.9
Aya-23 8 28.6 29.8 27.3 34.8 16.7 37.8 25.0 16.4 55.6 16.0 9.5 15.4 21.4 14.3 31.5 35.9 40.8 22.7 40.9
Llama 3.1 8 33.3 34.5 34.1 21.7 16.7 40.5 45.0 27.3 22.2 30.0 28.6 38.5 14.3 28.6 38.9 33.3 48.7 45.5 42.4
Llama 3 8 28.6 34.5 34.1 26.1 16.7 40.5 40.0 23.6 33.3 26.0 23.8 38.5 0.0 14.3 40.7 41.0 40.1 31.8 43.9
DeepSeek-R1 8 23.8 42.9 36.4 30.4 16.7 35.1 20.0 27.3 22.2 30.0 23.8 61.5 21.4 0.0 33.3 25.6 34.2 45.5 37.9
GLM 4 9 47.6 56.0 50.0 39.1 33.3 21.6 20.0 25.5 11.1 22.0 19.0 30.8 14.3 14.3 27.8 23.1 43.4 54.5 56.1
Gemma 2 9 42.9 52.4 20.5 56.5 33.3 27.0 20.0 27.3 11.1 28.0 9.5 46.2 21.4 57.1 20.4 17.9 36.8 40.9 42.4
Mistral-Nemo 12 28.6 44.0 34.1 17.4 16.7 24.3 30.0 30.9 22.2 30.0 28.6 23.1 35.7 42.9 29.6 15.4 40.8 27.3 36.4
StableLM 2 12 23.8 31.0 20.5 30.4 16.7 10.8 5.0 7.3 11.1 8.0 9.5 0.0 7.1 14.3 24.1 10.3 32.2 27.3 24.2
Llama 2 13 23.8 26.2 6.8 13.0 33.3 16.2 5.0 14.5 11.1 14.0 0.0 30.8 7.1 0.0 16.7 7.7 13.2 13.6 15.2
Phi 3 14 19.0 38.1 34.1 30.4 16.7 40.5 50.0 32.7 33.3 30.0 28.6 30.8 14.3 42.9 38.9 38.5 40.8 29.5 39.4
Medium-sized Models $(>15B\&\leq 200B)$
Mistral Small 22 0.0 13.1 0.0 4.3 0.0 0.0 0.0 1.8 0.0 2.0 4.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.5
Gemma 2 27 4.8 15.5 6.8 13.0 0.0 13.5 15.0 5.5 11.1 6.0 4.8 7.7 7.1 0.0 9.3 12.8 3.3 6.8 0.0
Aya-23 35 0.0 11.9 9.1 0.0 0.0 2.7 0.0 3.6 0.0 4.0 0.0 15.4 0.0 0.0 1.9 0.0 0.0 4.5 4.5
Command-R 35 0.0 6.0 4.5 4.3 0.0 2.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.9 2.6 0.7 4.5 0.0
Alfred 40 0.0 1.2 2.3 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 5.6 0.0 0.7 2.3 0.0
Mixtral 8x7 0.0 1.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.9 0.0 0.0 0.0 0.0
Llama 2 70 0.0 0.0 0.0 0.0 0.0 2.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.9 0.0 0.0 2.3 0.0
Llama 3 70 0.0 26.2 15.9 17.4 0.0 21.6 10.0 9.1 0.0 10.0 0.0 15.4 7.1 14.3 13.0 10.3 0.7 9.1 1.5
Llama 3.1 70 4.8 31.0 22.7 17.4 0.0 13.5 10.0 12.7 11.1 14.0 4.8 23.1 7.1 28.6 16.7 12.8 2.0 6.8 6.1
DeepSeek-R1 70 4.8 26.2 18.2 8.7 0.0 21.6 5.0 10.9 0.0 12.0 9.5 30.8 0.0 0.0 31.5 15.4 0.7 4.5 0.0
Qwen 2 72 0.0 13.1 11.4 8.7 0.0 5.4 15.0 7.3 0.0 8.0 4.8 7.7 14.3 28.6 5.6 10.3 2.6 11.4 3.0
Command-R+ 104 4.8 16.7 9.1 17.4 16.7 0.0 0.0 7.3 0.0 4.0 0.0 7.7 7.1 0.0 1.9 0.0 3.3 4.5 1.5
Mistral-Large 123 4.8 23.8 18.2 13.0 0.0 2.7 5.0 7.3 0.0 8.0 4.8 15.4 0.0 14.3 9.3 2.6 0.7 4.5 1.5
Mixtral 8x22 0.0 3.6 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 3.7 2.6 0.0 0.0 1.5
WizardLM 2 8x22 0.0 13.1 2.3 0.0 0.0 5.4 5.0 3.6 0.0 4.0 0.0 15.4 0.0 0.0 5.6 5.1 0.0 4.5 1.5
Large Models $(>200B)$
Llama 3.1 405 42.9 67.9 68.2 47.8 33.3 37.8 40.0 41.8 22.2 38.0 33.3 53.8 28.6 14.3 55.6 33.3 55.9 63.6 59.1
DeepSeek-R1 671 81.0 94.0 93.2 91.3 100.0 70.3 65.0 69.1 55.6 70.0 57.1 84.6 71.4 100.0 81.5 64.1 79.6 86.4 84.8
ChatGPT 3.5 unk 71.4 91.7 90.9 100.0 50.0 73.0 60.0 74.5 44.4 76.0 71.4 92.3 71.4 71.4 75.9 69.2 85.5 88.6 83.3
Claude 3.5 S unk 42.9 85.7 88.6 100.0 50.0 70.3 75.0 89.1 66.7 90.0 81.0 100.0 92.9 85.7 85.2 69.2 88.2 86.4 86.4
ChatGPT 4o unk 52.4 86.9 90.9 95.7 66.7 78.4 60.0 74.5 55.6 76.0 66.7 92.3 78.6 85.7 74.1 64.1 83.6 86.4 83.3
ChatGPT 4o m unk 38.1 79.8 68.2 87.0 16.7 56.8 65.0 52.7 22.2 54.0 47.6 76.9 42.9 71.4 64.8 51.3 61.8 86.4 81.8
Gemini 2.0 F unk 33.3 81.0 84.1 87.0 16.7 70.3 70.0 72.7 66.7 76.0 66.7 84.6 71.4 85.7 68.5 53.8 80.9 86.4 84.8
LeChat unk 38.1 65.5 61.4 47.8 50.0 24.3 50.0 34.5 44.4 34.0 28.6 38.5 35.7 42.9 35.2 35.9 39.5 40.9 47.0
Table 9: Complete results of experiments on fine-grained categories
G Fine-grained Category Description
Table 10 contains explanations and statistics for the fine-grained categories we created for the exam dataset.
Category name Category Description #
V_idiom a question relating to idioms or metaphors 21
V_noun a vocabulary question relating to a noun 84
V_verb a vocabulary question relating to a verb 44
V_adjective a vocabulary question relating to an adjective 23
V_native a vocabulary question where the correct answer is a word exclusive to Luxembourgish 6
G_gender a question relating to determining the gender of a determinant, adjective, or pronoun 37
G_pronouns a question relating to pronouns 20
G_verbs a question relating to conjugations of verbs 55
G_adjective a question relating to adjectives 9
G_tense a question relating to the correct use of tenses 50
G_past a question relating to the past tense 21
G_present a question relating to the present tense 13
G_subjunctive a question relating to the subjunctive tense 14
G_modal a question relating to modal verbs 7
G_preposition a question relating to the correct use of prepositions or conjunctions 54
G_cases a question relating to nominative,genitive,dative, or accusative 39
RC_TF a reading comprehension question where the answer is either true or false 152
RC_Simple a reading comprehension question that does not require to understand the entire text excerpt to answer correctly 72
RC_Complex a reading comprehension question requiring a deeper understanding of the entire text excerpt to answer correctly 86
Table 10: Statistics for fine-grained categories for exam questions.
H Fine-grained Category Annotation Prompt
We use the following prompt to categorise the exam questions:
I have a language test for Luxembourgish where I would like you to categorise every question
into different types of skill they are testing. For every question and every category,
determine and mark whether or not a given question belongs to a given category, as follows:
0, the question does not belong to the category
1, the question belongs to the category
Write the output in tabular form like TSV.
The file I am uploading are some questions that I already manually annotated.
Use it as a guideline.
The list of categories and explanations is as follows:
vocabulary: any question relating to vocabulary
grammar: any question relating to grammar
vocabulary_idiom: a question relating to idioms or metaphors
vocabulary_noun: a vocabulary question relating to a noun
vocabulary_verb: a vocabulary question relating to a verb
vocabulary_adjective: a vocabulary question relating to an adjective
vocabulary_native_word: a vocabulary question where the correct answer is a
word exclusive to Luxembourgish
grammar_gender: a question relating to determining the gender of a
determinant, adjective, or pronoun
grammar_pronouns: a question relating to pronouns
grammar_verbs:  a question relating to conjugations of verbs
grammar_adjectives:  a question relating to adjectives
grammar_tense: a question relating to the correct use of tenses
grammar_tense_past: a question relating to the past tense
grammar_tense_present: a question relating to the present tense
grammar_tense_subjunctive: a question relating to the subjunctive tense
grammar_verbs_modal_verbs: a question relating to modal verbs
grammar_preposition_conjunction: a question relating to the correct use
of prepositions or conjunctions
grammar_cases: a question relating to nominative,genitive,dative, or accusative
reading_comprehension_true_false: a reading comprehension question
where the answer is either true or false
reading_comprehension_simple: a reading comprehension question that
does not require to understand the entire text excerpt to answer correctly
reading_comprehension_complex: a reading comprehension question that
requires a deeper understanding of the entire text excerpt to answer correctly
I LuxGen Evaluation Results
Table 11 shows the full evaluation results of the LuxGen experiments over three runs as well as the average performances.
Small Description News Headline
Model Size Exam Score Grammar Adequacy Grammar Adequacy
Run 1 Run 2 Run 3 Average Run 1 Run 2 Run 3 Average Run 1 Run 2 Run 3 Average Run 1 Run 2 Run 3 Average
Small Models $(\leq 15B)$
Qwen 2 0.5 0.220 0.174 0.318 0.290 0.261 0.451 0.340 0.330 0.374 0.124 0.160 0.352 0.212 0.303 0.303 0.433 0.347
Llama 3.2 1 0.044 0.534 0.520 0.582 0.545 0.521 0.572 0.550 0.548 0.338 0.256 0.418 0.337 0.228 0.257 0.525 0.337
Qwen 2 1.5 0.286 0.538 0.433 0.502 0.491 0.553 0.462 0.442 0.485 0.226 0.162 0.406 0.265 0.398 0.377 0.492 0.422
StableLM 2 1.6 0.094 0.025 0.232 0.332 0.196 0.284 0.298 0.212 0.265 0.214 0.256 0.390 0.287 0.360 0.285 0.513 0.386
Gemma 2 2 0.278 0.456 0.642 0.568 0.555 0.533 0.535 0.658 0.575 0.304 0.408 0.438 0.383 0.608 0.492 0.495 0.532
Llama 3.2 3 0.321 0.830 0.782 0.812 0.808 0.689 0.737 0.648 0.691 0.692 0.702 0.656 0.683 0.653 0.647 0.695 0.665
Phi 3 3.8 0.300 0.045 0.362 0.477 0.294 0.391 0.320 0.362 0.358 0.220 0.202 0.434 0.285 0.327 0.320 0.600 0.416
Phi 3.5 3.8 0.101 0.267 0.258 0.287 0.271 0.243 0.150 0.147 0.180 0.118 0.065 0.488 0.224 0.240 0.328 0.625 0.398
Qwen 2 7 0.373 0.456 0.725 0.687 0.623 0.644 0.660 0.688 0.664 0.380 0.336 0.546 0.421 0.578 0.597 0.548 0.574
Llama 2 7 0.141 0.054 0.275 0.238 0.189 0.296 0.275 0.225 0.265 0.146 0.180 0.424 0.250 0.425 0.408 0.522 0.452
WizardLM 2 7 0.267 0.140 0.440 0.585 0.388 0.529 0.572 0.432 0.511 0.298 0.248 0.466 0.337 0.505 0.553 0.615 0.558
Mistral 7 0.325 0.242 0.767 0.650 0.553 0.634 0.695 0.680 0.670 0.518 0.474 0.548 0.513 0.548 0.537 0.633 0.573
Aya-23 8 0.322 0.377 0.470 0.708 0.518 0.510 0.515 0.448 0.491 0.086 0.142 0.446 0.225 0.340 0.372 0.557 0.423
Llama 3.1 8 0.384 0.655 0.658 0.755 0.689 0.610 0.487 0.620 0.572 0.684 0.644 0.704 0.677 0.528 0.643 0.603 0.592
Llama 3 8 0.357 0.624 0.690 0.722 0.678 0.630 0.643 0.627 0.633 0.322 0.310 0.554 0.395 0.600 0.578 0.630 0.603
DeepSeek-R1 8 0.338 0.628 0.592 0.667 0.629 0.533 0.507 0.482 0.507 0.576 0.458 0.532 0.522 0.637 0.592 0.543 0.591
GLM 4 9 0.421 0.721 0.783 0.795 0.766 0.591 0.590 0.715 0.632 0.322 0.252 0.402 0.325 0.580 0.658 0.578 0.606
Gemma 2 9 0.360 0.622 0.807 0.818 0.749 0.623 0.645 0.762 0.677 0.738 0.584 0.652 0.658 0.730 0.678 0.550 0.653
Mistral-Nemo 12 0.340 0.698 0.737 0.753 0.729 0.663 0.718 0.698 0.693 0.578 0.502 0.572 0.551 0.707 0.752 0.677 0.712
StableLM 2 12 0.235 0.263 0.577 0.348 0.396 0.437 0.320 0.325 0.361 0.202 0.190 0.442 0.278 0.323 0.432 0.572 0.442
Llama 2 13 0.147 0.140 0.438 0.315 0.298 0.286 0.230 0.295 0.270 0.200 0.236 0.512 0.316 0.460 0.468 0.583 0.504
Phi 3 14 0.370 0.124 0.358 0.435 0.306 0.331 0.328 0.242 0.300 0.212 0.152 0.422 0.262 0.307 0.263 0.582 0.384
Medium-sized Models $(>15B\&\leq 200B)$
Mistral Small 22 0.024 0.675 0.843 0.848 0.789 0.633 0.620 0.692 0.648 0.812 0.772 0.738 0.774 0.750 0.780 0.620 0.717
Gemma 2 27 0.066 0.648 0.853 0.882 0.794 0.663 0.728 0.808 0.733 0.776 0.674 0.730 0.727 0.780 0.760 0.607 0.716
Aya-23 35 0.036 0.711 0.620 0.667 0.666 0.631 0.697 0.573 0.634 0.450 0.374 0.508 0.444 0.558 0.558 0.625 0.581
Command-R 35 0.021 0.517 0.653 0.745 0.638 0.621 0.658 0.667 0.649 0.182 0.423 0.642 0.416 0.617 0.688 0.670 0.658
Alfred 40 0.011 0.206 0.402 0.503 0.370 0.381 0.420 0.405 0.402 0.272 0.350 0.272 0.298 0.495 0.573 0.342 0.470
Mixtral 8x7 0.003 0.226 0.470 0.503 0.400 0.454 0.437 0.512 0.468 0.292 0.258 0.504 0.351 0.630 0.638 0.663 0.644
Llama 2 70 0.005 0.464 0.632 0.592 0.562 0.599 0.505 0.623 0.576 0.152 0.212 0.482 0.282 0.432 0.495 0.622 0.516
Llama 3 70 0.101 0.788 0.815 0.860 0.821 0.719 0.778 0.763 0.753 0.336 0.294 0.580 0.403 0.735 0.722 0.718 0.725
Llama 3.1 70 0.121 0.842 0.865 0.895 0.867 0.731 0.818 0.800 0.783 0.712 0.730 0.810 0.751 0.777 0.800 0.700 0.759
DeepSeek-R1 70 0.107 0.872 0.865 0.838 0.858 0.597 0.767 0.652 0.672 0.772 0.630 0.724 0.709 0.762 0.785 0.618 0.722
Qwen 2 72 0.068 0.776 0.803 0.708 0.762 0.684 0.652 0.633 0.656 0.584 0.516 0.730 0.610 0.747 0.765 0.655 0.722
Command-R+ 104 0.057 0.705 0.688 0.818 0.737 0.627 0.683 0.700 0.670 0.646 0.682 0.754 0.694 0.697 0.777 0.693 0.722
Mistral-Large 123 0.076 0.752 0.777 0.832 0.787 0.663 0.682 0.752 0.699 0.852 0.790 0.822 0.821 0.827 0.818 0.743 0.796
Mixtral 8x22 0.011 0.515 0.622 0.790 0.642 0.566 0.567 0.528 0.554 0.620 0.604 0.700 0.641 0.780 0.782 0.700 0.754
WizardLM 2 8x22 0.037 0.216 0.520 0.428 0.388 0.407 0.360 0.392 0.386 0.404 0.390 0.600 0.465 0.720 0.740 0.658 0.706
Large Models $(>200B)$
Llama 3.1 405 0.552 0.776 0.762 0.755 0.764 0.666 0.695 0.623 0.661 0.874 0.814 0.854 0.847 0.820 0.803 0.742 0.788
DeepSeek-R1 671 0.804 0.980 0.857 0.890 0.909 0.630 0.705 0.760 0.698 0.866 0.658 0.653 0.726 0.800 0.817 0.647 0.754
ChatGPT 3.5 unk 0.809 0.632 0.833 0.750 0.738 0.659 0.718 0.682 0.686 0.610 0.716 0.614 0.647 0.595 0.670 0.643 0.636
Claude 3.5 Sonnet unk 0.834 0.874 0.752 0.875 0.834 0.739 0.667 0.780 0.728 0.884 0.856 0.740 0.827 0.862 0.860 0.777 0.833
ChatGPT 4o unk 0.795 0.873 0.902 0.910 0.895 0.726 0.673 0.760 0.720 0.728 0.834 0.720 0.761 0.833 0.845 0.765 0.814
ChatGPT 4o mini unk 0.666 0.863 0.735 0.858 0.819 0.500 0.608 0.680 0.596 0.766 0.754 0.684 0.735 0.705 0.718 0.720 0.714
Gemini 2.0 Flash unk 0.768 0.814 0.792 0.815 0.807 0.606 0.555 0.512 0.557 0.544 0.550 0.512 0.535 0.698 0.777 0.675 0.717
LeChat unk 0.438 0.639 0.505 0.543 0.562 0.444 0.630 0.597 0.557 0.778 0.578 0.455 0.604 0.705 0.725 0.597 0.676
Table 11: Detailed LuxGen Evaluation Results